Fermi National Laboratory
 

Operation Summary for December 16, 2009 - December 18, 2009

 

Wednesday Dec. 16
The day shift began with Operations monitoring the Tevatron (TeV) store 7440 (85.49E30) and the Recycler stash (185.2E10), with the Antiproton Source stacking (11.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, at 12:45 PM, the Pbar Debuncher extraction kicker power supply (D:EKIK) faulted. EE Support personnel completed its repair and Pbar resumed stacking by the end of the shift.

During the evening shift, Operations terminated store 7440 at 9:30 PM. Operations put the TeV into HEP shot setup and established store 7441 at 11 PM with an average initial luminosity (AIL) of 266.0E30. Operations disabled the beam to MTest at 7 PM.

Thursday Dec. 17
The midnight shift began with Operations monitoring store 7441 (206.89E30) and the stash (66.4E10), with Pbar stacking (29.2E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 4:35 AM.

During the day shift, Main Injector RF problems caused beam inefficiencies and problems with coalescing, experts were able to solve the problem by the end of the shift.

During the evening shift, Operations terminated store 7441 at 4:04 PM. Operations put the TeV into HEP shot setup and established store 7442 at 5:45 PM with an average initial luminosity (AIL) of 275.0E30.

Friday Dec. 18
The midnight shift began with Operations monitoring store 7442 (105.69E30) and the stash (210.0E10), with Pbar stacking (19.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 4:05 AM. Store 7442 was terminated at 8:04 AM.

The Plans for Friday and the Weekend
The plan for today is to allow D0 an 8 to 10 hour access to repair a bad preamp. Other accesses will be made to take advantage of this down time. Shot setup will follow.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/18/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory